A swimmer runs horizontally off a diving board with a speed of 3.14 m/s and hits the water a horizontal distance of 2.16 m from the end of the board. How high above the water was the diving board?
Answer #1
How long did it take to hit the water?

time= 2.16/3.14 seconds

hf=ho- 1/2 g t^2
0=ho-4.9 t^2 solve for ho,knowing t